OpenStudy (anonymous):

Lenny runs 100-meters course in 25 seconds. Gary runs a 450-meter course in 112.5 seconds. Bruford runs a 950-meter course in 237.5 seconds. Which athlete is the fastest? Round each speed to one decimal place.

OpenStudy (anonymous):

\[\left\{\frac{100}{25},\frac{450}{112},\frac{950}{237.5}\right\} \]\[\{4.,4.01786,4.\} \]